# Uncovering Hidden Connections: Iterative Tracking and Reasoning for Video-grounded Dialog

### v1

@@ 4.1 Dataset
In evaluating our model, we used a benchmark dataset tailored for video-grounded dialog. Sourced from the Charades human-activity dataset [46], it captures authentic human interactions about video snippets. Two notable versions of this dataset exist: AVSD@DSTC7 [31] and AVSD@DSTC8 [32]. It contains 7,659 training dialogs, 1,787 for validation, and 1,710 each for DSTC7 and DSTC8 testing in the text generation task. Although both testing sets have the same number of dialogs, more nuanced utterance comprehension as well as long-term history dependency capture is more critical in DSTC8 due to longer sentences. Each dialog, tied to a video, contains ten paired question-response utterances. Dataset specifics are detailed in Table II.
TABLE III: Performance comparison with the latest methods on the AVSD@DSTC8 dataset. The best results are highlighted in bold and the second in underline.
@@
ITR - 0.762 0.641 0.543 0.460 0.298 0.607 1.285
@@ 4.2.1 Implementation Details
We trained our model for 20 epochs using the Adam optimizer. The learning rate was set at 4.5e-4 and reduced tenfold between the 5-th and 10-th epochs. Configurations included a frame count ($F$) of 15, object number ($O$) of 6, and attention heads ($J$) of 3. We set the selection threshold ($p$) at 0.6 and iterated the reasoning network three times. The feature dimension ($d$) was 768, and batches were organized in sets of 32. During decoding, we employed a beam search with a width of 5, a sequence length capped at 20, and a length penalty of 0.2. Hyperparameters were optimized using a grid search on the validation set. Experiments were conducted using PyTorch on a single NVIDIA A100-PCIE-40GB GPU with CUDA 11.0.
@@ 4.3 Performance Comparison
We evaluated the performance of our proposed framework against several state-of-the-art benchmarks. This comparison includes generalized methodologies (e.g., MTN [11], BiST [18], SCGA [14]), as well as GPT-based strategies (e.g., PDC [15], RLM [12], T5RLM [23]) on the AVSD@DSTC7 and AVSD@DSTC8 benchmark datasets. The results for the AVSD@DSTC7 dataset are summarized in Table I, with the following notable insights:
  • Our model exhibited substantial improvements over earlier baselines such as HMA, RMFF, and EE-DMN, with metrics like CIDEr showing a boost from 0.856 (HMA) to 1.331 (Our model), highlighting the distinct advantages of our model.
  • Models leveraging pretrained language constructs like GPT-2 (PDC, RLM, T5RLM, and JVIT) generally outperformed non-pretrained counterparts, confirming the efficacy of pretrained models in text generation.
  • Our model surpassed all compared methods, achieving the highest scores across almost all evaluation metrics. This demonstrates its effective handling of the text generation task, with results such as a BLEU-4 score of 0.469, a 3.5% improvement over the next-best JVIT model.
The performance of our proposed framework, juxtaposed against several state-of-the-art methods on the AVSD@DSTC8 dataset, is delineated in Table III. Key observations and insights drawn from the comparative results are as follows:
  • Across most methods, there is a noticeable decrease in performance from AVSD@DSTC7 to AVSD@DSTC8, evidenced by metrics like BLEU-4, METEOR, ROUGE-L, and CIDEr. For example, the BLEU-4 score of RLM drops from 0.459 to 0.445. Similar trends are seen in models like MTN-P and SCGA. This drop might be attributed to differences in the datasets rather than solely the performance of the models.
  • Our proposed model continues to excel, showing improvements across various evaluation metrics for the AVSD@DSTC8 benchmark. Specifically, our method outperforms DialogMCF by 2.4% and 2.5% for BLEU-4 and CIDEr, respectively. These results emphasize the robustness and capability of our designed framework.
In conclusion, our methodology maintains a consistent edge over other contenders on both AVSD@DSTC7 and AVSD@DSTC8 datasets. This dominance, mirrored across diverse evaluation metrics, accentuates the dexterity and resilience of our devised approach. In particular, the repeated attainment of pinnacle scores across a myriad of metrics underscores the holistic efficacy of our model.
TABLE IV: Performance comparison among the variants of our model on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
@@
TABLE V: Performance comparison among the variants of different generation methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
@@
4.4 Ablation Study
The goal of our ablation study is to critically evaluate the individual contribution of distinct modules within our ITR model. By strategically removing certain components, we can assess their significance in the holistic performance of our model. We designed four derivatives to accomplish this:
  • w/o visual encoder. In this configuration, the proposed visual encoder was omitted, and the initial feature $\mathbf{E}_{v}$ and question feature $\mathbf{E}_{q}^{txt}$ emanating from the textual encoder were directly inputted into the generator.
  • w/o GPT-2 decoder. In this derivative, the GPT-2 decoder was supplanted with the LSTM-based auto-regressive decoder that was previously utilized in [11].
Table IV presents a meticulous evaluation of the ITR model against its various derivatives using the AVSD@DSTC7 dataset. The empirical findings from this analysis solidify the effectiveness of the ITR model, accentuating the significance of each constituent component within the architecture. We glean the following salient insights from the data:
  • The w/o textual encoder variant, which omits the textual encoder, underperforms in comparison to the full-fledged ITR model. This underscores the criticality of dialog history in grasping the nuances and intent behind user queries.
  • The model that bypasses the gate mechanism (w/o gate mechanism) delivers suboptimal results compared to the complete ITR. This result reinforces the efficiency of the gate mechanism in fusing cross-modal information.
Overall, this ablation study clearly demarcates the contribution of each module to the performance of our model, highlighting the robust design of the proposed ITR framework.
4.5 Decoding Methods
Table V furnishes an in-depth analysis of various decoding strategies deployed within the domain of multimodal dialog generation. A synopsis of our findings is delineated below:
  • Greedy Search. While Greedy search that selects the most probable subsequent word in each instance exhibits laudable performance (BLEU-1 score of 0.756), its methodical nature can sometimes overlook globally coherent or intricate responses due to its short-sighted selection strategy.
  • Nucleus Sampling. Posting a BLEU-1 score of 0.695, nucleus sampling amalgamates randomness by selecting from a subset of the most likely subsequent words. The infusion of diversity within responses, while a strength, might underpin its subdued coherence in comparison to its counterparts within video-anchored dialogs.
  • Contrastive Search. Registering a commendable BLEU-1 score of 0.773, the contrastive modus operandi evaluates divergent options, possibly assisting in sifting through the subtleties of video-grounded retorts. Notwithstanding its edge over the first two, it could not surpass beam search in our assessment.
  • Beam Search. Clinching the pinnacle with a BLEU-1 score of 0.782, beam search meticulously appraises multiple trajectories during decoding, culminating in the selection of the most contextually rich sequence. Its supremacy across the board testifies to its aptness for tasks necessitating a precise grounding in stimuli like videos.
A salient feature of video-grounded dialog is its deterministic response modality, anchored firmly in the video and its captions. This deterministic nature diverges from the more fluid and open-ended nature of general open-domain dialogs. It’s evident, in light of this characteristic, that the beam search mechanism emerges as particularly adept for decoding tasks within this milieu.
@@ 4.5 Decoding Methods
(c) The number of iterations.
Figure 4: Parameter analysis of our model in terms of the (a) the number of dialog turns, (b) the selection threshold $p$, (c) the number of iterations.
@@ 4.6 Parameter Analysis
In this section, we delved deeper into the repercussions of three principal hyperparameters: the number of dialog turns, the selection threshold $p$, and the number of iterations in the ITR model.
@@ 4.6.1 On the number of dialog turns
Dialog turns in our model encapsulate the iterative interaction history, serving as a critical backdrop against which current interactions are contextualized. We embarked on a quest to discern the optimal number of dialog turns that our model should consider for peak performance.
As showcased in Figure 4(a), our experiments spanned dialog histories ranging from 1 to 10 turns. A discernible trend emerged: both the BLEU-4 and CIDEr metrics initially surged, peaking at a dialog history length of 4, before exhibiting a downturn. This oscillation underscores a salient feature: while history is undeniably beneficial, inundating the model with an overabundance of it seems counterproductive. The decline in performance with longer dialog histories hints at the introduction of noise or redundant information, which potentially obfuscates more pertinent recent interactions.
@@ 4.6.2 On the selection threshold $p$
The selection threshold $p$ plays a pivotal role in filtering and modulating the information processed by the ITR model. To pinpoint its optimal value, we scrutinized the performance of our model across varying threshold levels.
The outcomes, illustrated in Figure 4(b), reveal a distinct pattern. The CIDEr metric, along with the BLUE-4 score, starts off with a surge, culminating in a peak at a threshold of 0.6, post which there is a decline. This showcases that extreme values on either end of the spectrum aren’t conducive. While an extremely stringent threshold might deprive the model of necessary data, an overly lenient one might flood it with superfluous information.
4.6.3 On the number of iterations
@@
Our data, highlighted in Figure 4(c), offers an answer. Both BLEU-4 and CIDEr metrics register peak performances at three iterations, suggesting that this is the sweet spot. A deeper dive reveals that beyond this point, adding more iterations seems counterproductive. This could be attributed to the additional computational overhead and potential for overfitting that come with more iterations.
@@ 4.7 Case Study
Deep-diving into the applied implications of our ITR model, we present a series of case studies sourced from the AVSD@DSTC7 dataset. While quantitative evaluations provide insights into the overall performance of our model, case studies offer a qualitative lens to appreciate its nuanced behaviors and decision-making rationale. In Figure 5, we offer a visual manifestation of two exemplar cases: Case A and Case B. Each case elucidates the action of our model by illustrating the subject path and object path that the model opts to traverse en route to deriving the final response.
In Case A, our model demonstrates its competence in sifting through the dialog history, isolating the information salient to the current interaction. It illustrates how the model constructs a subject path that accurately reflects the central theme of the inquiry, as well as an object path that pinpoints the specifics the user is inquiring about. The concordance between these paths and the final response manifests the coherence and logic embedded within the operations of the model. The correct answer is derived by linking the information from the paths, showcasing the intelligent decision-making of our model. Moreover, the traceable path structure affirms the interpretability of the ITR model. This fosters a level of trust, as the operations within the model can be tracked and understood rather than being concealed as a “black box”.
@@ 5 Conclusion And Future Work
In this paper, we introduce the Iterative Tracking and Reasoning (ITR) framework tailored for video-grounded dialog. The ITR framework comprises three main components: a textual encoder, a visual encoder, and a generator. Our textual encoder is equipped with a path tracking and aggregation strategy, which meticulously aggregates context pertinent to the user’s question, thus augmenting textual comprehension. In the visual domain, we introduce a multimodal iterative reasoning network within our visual encoder. This network delves deep into visual semantics that resonate with the user’s inquiry, thereby bolstering visual understanding. Merging these enriched insights, we employ the pre-trained GPT-2 model as our response generator to craft the final response. Rigorous evaluations across two benchmark datasets vouch for the preeminence of our ITR model over prevailing techniques.
In the future, we intend to explore multimodal large language models, facilitating research in the field of video-grounded dialog. Besides, we also plan to study video-grounded dialog in more perspectives, e.g., egocentric video dialog.

### v2

@@ 4.1 Dataset
We employed three publicly available datasets to evaluate our model, including AVSD@DSTC7 [34], AVSD@DSTC8 [35], and VSTAR [49]. The first two datasets are both derived from the Charades human-activity dataset [50] that captures authentic human interactions about video clips. The training and validation sets in both datasets are the same, while the testing set is completely different. Due to longer sentences in DSTC8, more nuanced utterance comprehension as well as long-term history dependency capture is more critical. Each dialog, tied to a video, contains ten paired question-answer utterances. VSTAR is a large-scale video-grounded dialog understanding dataset, which is collected from 395 TV series (185K 90-second clips) with carefully cleaned dialogs and metadata information. The sheer volume of data exacerbates the complexity of this dataset. Dataset specifics are detailed in Table II.
TABLE III: Performance comparison with the latest methods on the AVSD@DSTC8 dataset. The best results are highlighted in bold and the second is underlined.
@@
ISR - 0.762 0.641 0.543 0.460 0.298 0.607 1.285
TABLE IV: Performance comparison with the latest methods on the VSTAR dataset. The best results are highlighted in bold and the second is underlined.
Method BLEU-4 METEOR ROUGE-L CIDEr
OpenViDial coarse [52] 0.006 0.035 0.063 0.066
RLM [15] 0.010 0.032 0.061 0.079
AVDT [49] 0.014 0.040 0.080 0.137
ISR 0.020 0.044 0.081 0.150
@@ 4.2.1 Implementation Details
For model architecture, we fine-tuned the GPT-2 generator and froze the Faster R-CNN model as well as skipping the pre-trained text embedding model. We trained our model for 20 epochs using the Adam optimizer. The learning rate was set at 4.5e-4 and reduced tenfold between the 5-th and 10-th epochs. Configurations included a frame count ($F$) of 15, object number ($O$) of 6, and attention heads ($J$) of 3. The feature dimension ($d$) was 768, and batches were organized in sets of 32. During decoding, we employed a beam search with a width of 5, a sequence length capped at 20, and a length penalty of 0.2. Hyperparameters, the selection threshold ($p$) and iteration number ($I$), were separately set to 0.6 and 3 using a grid search on the validation set, where the range of $p$ is {0.0, 0.2, 0.4, 0.6, 0.8, 1.0} and $I$ ranges from 0 to 5. Experiments were conducted using PyTorch on a single NVIDIA A100-PCIE-40GB GPU with CUDA 11.0.
@@ 4.3 Performance Comparison
We evaluated the performance of our proposed framework against several state-of-the-art baselines. This comparison includes generalized methods (e.g., MTN [14], BiST [21], SCGA [17]), as well as GPT-based ones (e.g., PDC [18], RLM [15], T5RLM [26]) on the AVSD@DSTC7 and AVSD@DSTC8 datasets. The results for the AVSD@DSTC7 dataset are summarized in Table I, with the following notable insights:
  • Models leveraging pretrained language constructs like GPT-2 (PDC, RLM, T5RLM, and JVIT) generally outperformed non-pretrained counterparts, confirming the efficacy of pretrained models in text generation. This also provides a basis for subsequent work using pre-trained model.
  • Our model surpassed all compared methods, achieving the highest scores across almost all evaluation metrics. For example, ISR achieves a 3.5% relative gain in BLEU-4 score compared to the JVIT baseline. These results not only validate the effectiveness of our overall framework but also reveal the necessity of mining deeper clues for question understanding.
The performance of our proposed framework, juxtaposed against several state-of-the-art methods on the AVSD@DSTC8 dataset, is delineated in Table III. Key observations drawn from the comparative results are as follows:
  • Across most methods, there is a noticeable decrease in performance from AVSD@DSTC7 to AVSD@DSTC8, evidenced by metrics like BLEU-4, METEOR, ROUGE-L, and CIDEr. For example, the BLEU-4 score of RLM drops from 0.459 to 0.445. The observed decline in performance may be due to the disparities between the two datasets, particularly because the questions in the AVSD@DSTC8 testing set exhibit greater complexity.
  • Our proposed model demonstrates significant advancements, achieving improvements across multiple evaluation metrics on the AVSD@DSTC8 benchmark. Notably, it surpasses DialogMCF by 2.4% in BLEU-4 and by 2.5% in CIDEr scores. These findings underscore the robustness and enhanced effectiveness of our model.
The evaluation results of our ISR model on the VSTAR dataset are shown in Table IV. And we can derive the following conclusions from this:
  • The performance of all methods on this dataset is relatively poor. We believe that this is on the one hand caused by the large size of the data, while on the other hand, TV serials are perhaps more difficult for models to understand compared to real videos.
  • On this extremely difficult dataset, our model also obtains optimal performance compared to the latest baselines. This phenomenon again illustrates the effectiveness of our path search and aggregation strategy and iterative reasoning network for mining the hidden semantics of complex data.
In conclusion, our methodology maintains a consistent edge over other contenders on AVSD@DSTC7, AVSD@DSTC8, and VSTAR datasets. This dominance, mirrored across diverse evaluation metrics, accentuates the dexterity and resilience of our devised approach. In particular, the consistent achievement of top scores across numerous metrics underscores the comprehensive effectiveness of our model for uncovering hidden connections.
TABLE V: Performance comparison among the variants of our model on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
@@
TABLE VI: Performance comparison with the large video-language models. The version of ChatGPT is “gpt-3.5-turbo” and the best performance is highlighted in bold.
Methods Objective Metric ChatGPT
BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr Accuracy Score
AVSD@DSTC7
Video-LLaMA-7B 0.360 0.244 0.165 0.115 0.210 0.367 0.383 28.533 2.551
Video-LLaVA-7B 0.422 0.291 0.200 0.142 0.239 0.418 0.447 33.158 2.918
ISR 0.782 0.655 0.552 0.469 0.305 0.619 1.331 51.843 3.630
AVSD@DSTC8
Video-LLaMA-7B 0.345 0.231 0.158 0.110 0.201 0.364 0.444 28.737 2.575
Video-LLaVA-7B 0.410 0.280 0.194 0.137 0.228 0.416 0.527 33.275 2.998
ISR 0.762 0.641 0.543 0.460 0.298 0.607 1.285 51.012 3.588
4.4 Ablation Study
The goal of our ablation study is to critically evaluate the individual contribution of distinct modules within our ISR model. By strategically removing certain components, we can assess their significance in the holistic performance of our model. We designed four derivatives to accomplish this:
  • w/o visual encoder. In this configuration, the proposed visual encoder is omitted, and the initial feature $\mathbf{E}_{v}$ and question feature $\mathbf{E}_{q}^{txt}$ emanating from the textual encoder are directly inputted into the generator.
  • w/o GPT-2 decoder. In this derivative, the GPT-2 decoder is supplanted with the LSTM-based auto-regressive decoder that is previously utilized in [14].
Table V presents a meticulous evaluation of the ISR model against its various derivatives using the AVSD@DSTC7 dataset. We can obtain the following salient insights from the results:
  • The w/o textual encoder variant, which omits the textual encoder, underperforms in comparison to the complete ISR model. This underscores the criticality of dialog history in grasping the nuances and intent behind user queries.
  • The model that bypasses the gate mechanism (w/o gate mechanism) delivers suboptimal results compared to the ISR. This result reinforces the effectiveness of the gate mechanism in fusing cross-modal information.
Overall, this ablation study demarcates the contribution of each module to the performance of our model, highlighting the robust design of the proposed ISR framework.
4.5 Comparison with LVLMs
It is well known that LVLMs have achieved excellent performance on several video-language tasks, e.g., video question answering and video captioning. Therefore, to verify their effectiveness on the video-grounded dialog task, we chose representative Video-LLaMA [37] and Video-LLaVA [53] methods to compare with ours, and the experimental results are shown in Table VI.
From the displayed results, we can see that our model significantly outperforms the two LVLMs over all evaluation metrics on two datasets. For a broad comparison, we have also utilized ChatGPT for evaluation [^8] https://github.com/mbzuai-oryx/Video-ChatGPT/tree/main/quantitative_evaluation., focusing on sentence semantics eliminates structure effects. It should be noted that Accuracy metric refers to whether the predicted answer semantically matches the ground truth, while the Score is an value between 0 and 5, with 5 indicating the highest meaningful match. In this case, our method also achieves the best performance. Crucially, compared to 7B parameters of the two LVLMs, our method yields such results using only about 550M parameters. These findings underscore the challenges involved in adapting LVLMs to video-grounded dialog task and further corroborate the efficacy of our method in uncovering the complex interplays within video-grounded dialog.
4.6 Decoding Methods
Table VII furnishes an in-depth analysis of various decoding strategies deployed within the domain of video-grounded dialog. A synopsis of our findings is delineated below:
  • Greedy Search. While greedy search that selects the most probable subsequent word in each instance exhibits laudable performance (BLEU-1 score of 0.756), it can sometimes overlook globally coherent or intricate responses due to its short-sighted selection strategy.
  • Nucleus Sampling. Posting a BLEU-1 score of 0.695, nucleus sampling amalgamates randomness by selecting from a subset of the most likely subsequent words. The introduction of diversity in responses might interfere with answer coherence in comparison to its counterparts within video-grounded dialogs.
  • Contrastive Search. Achieving a commendable BLEU-1 score of 0.773, the contrastive approach evaluates divergent options, potentially aiding in discerning the nuances of video-grounded responses. Despite its superiority over the first two methods, it could not surpass beam search in our assessment.
  • Beam Search. Clinching the pinnacle with a BLEU-1 score of 0.782, beam search meticulously evaluates multiple trajectories during decoding, ultimately selecting the most contextually rich sequence. Its dominance across all evaluations attests to its suitability for tasks requiring precise alignment with videos.
A notable feature of video-grounded dialog is its deterministic response modality, firmly anchored in the video and its caption. This deterministic nature differs from the more fluid and open-ended nature of general open-domain dialogs. It’s evident, in light of this characteristic, that the beam search mechanism emerges as particularly adept for decoding tasks within this situation.
TABLE VII: Performance comparison among the variants of different generation methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
@@ 4.6 Decoding Methods
(c) The number of iterations $I$.
Figure 4: Parameter analysis of our model in terms of the (a) the number of dialog turns, (b) the selection threshold $p$, (c) the number of iterations $I$.
@@ 4.7 Parameter Analysis
In this section, we explored in depth the implications of three key hyperparameters: the number of dialog turns, the selection threshold $p$, and the number of iterations $I$ in the ISR model.
@@ 4.7.1 On the number of dialog turns
Dialog turns in our model encapsulate the iterative interaction history, serving as a critical backdrop against which current interactions are contextualized. We undertook a quest to ascertain the ideal number of dialog turns for maximizing our model’s performance.
As showcased in Fig. 4(a), our experiments spanned dialog histories ranging from 1 to 10 turns. A discernible trend emerged: both the BLEU-4 and CIDEr metrics initially surged, peaking at a dialog history length of 4, before exhibiting a downturn. This oscillation underscores a salient feature: while history is undeniably beneficial, inundating the model with an overabundance of it seems counterproductive. The decline in performance with longer dialog histories hints at the introduction of noise or redundant information, which potentially obfuscates more pertinent recent interactions.
@@ 4.7.2 On the selection threshold $p$
The selection threshold $p$ plays a pivotal role in filtering and modulating the information processed by the ISR model. To pinpoint its optimal value, we scrutinized the performance of our model across varying threshold levels.
As illustrated in Fig. 4(b), the CIDEr metric, along with the BLUE-4 score, starts off with a surge, culminating in a peak at a threshold of 0.6, post which there is a decline. This showcases that extreme values on either end of the spectrum are not conducive. While an extremely stringent threshold might deprive the model of necessary data, an overly lenient one might flood it with superfluous information.
4.7.3 On the number of iterations $I$
@@
Our data, highlighted in Fig. 4(c), offers an answer. Both BLEU-4 and CIDEr metrics register peak performances at three iterations, suggesting that this is the sweet spot. A deeper examination reveals that beyond this point, adding more iterations seems counterproductive. This could be attributed to the additional computational overhead and potential for overfitting that come with more iterations.
@@ 4.8 Case Study
Delving deeply into the applied implications of our ISR model, we present a series of case studies sourced from the AVSD@DSTC7 dataset. While quantitative evaluations provide insights into the overall performance of our model, case studies offer a qualitative lens to appreciate its nuanced behaviors and decision-making rationale. In Fig. 5, we offer a visual presentation of two cases: Case A and Case B. Each case elucidates the action of our model by illustrating the subject path and object path that the model chooses to traverse en route to deriving the final response.
In Case A, our model demonstrates its competence in sifting through the dialog history, isolating the information salient to the current interaction. It illustrates how the model constructs a subject path that accurately reflects the central theme of the inquiry, as well as an object path that pinpoints the specifics the user is inquiring about. The concordance between these paths and the final response manifests the coherence and logic embedded within the operations of the model. The correct answer is derived by linking the information from the paths, showcasing the intelligent decision-making of our model. Moreover, the traceable path structure affirms the interpretability of the ISR model. This fosters a level of trust, as the operations within the model can be tracked and understood rather than being concealed as a “black box”.
@@ 5 Conclusion And Future Work
In this paper, we introduce the Iterative Search and Reasoning (ISR) framework tailored for video-grounded dialog. The ISR framework comprises three main components: a textual encoder, a visual encoder, and a generator. Our textual encoder is equipped with a path search and aggregation strategy, which meticulously aggregates context pertinent to the user’s question, thus augmenting textual comprehension. In the video domain, we design a multimodal iterative reasoning network within our visual encoder. This network delves deeply into visual semantics that resonates with the user’s inquiry, thereby bolstering visual understanding. Merging these enriched insights, we employ the pre-trained GPT-2 model as our response generator to produce the final response. Rigorous evaluations across three datasets attest to the superiority of our ISR model over existing techniques.
In the future, we intend to explore multimodal large language models, facilitating research in the field of video-grounded dialog. Besides, we also plan to study video-grounded dialog in more perspectives, e.g., egocentric video-grounded dialog.